# Quantifying Classifier Utility under Local Differential Privacy

### v2

@@ Abstract
1 Introduction
@@
Figure 1: From empirical data utility to theoretical data utility. This paper provides a theoretical quantification of classifier utility under LDP mechanisms by connecting the concentration analysis of LDP mechanisms with the robustness analysis of classifiers.
@@
Classifiers are functions that map input data to class labels, playing a fundamental role in industrial applications ranging from predictive modeling and data clustering to image recognition [38, 35, 30]. When deployed as services, classifiers require users to provide data inputs, which often contain sensitive information such as medical records or financial data, raising significant privacy concerns. While users wish to benefit from classification services, they may be reluctant to reveal their actual data. A common approach to address this concern is data perturbation, e.g. adding noise to numerical data or applying blur filters to images before sending them to the classifier. Data perturbation remains a lightweight and intuitive solution for privacy-preserving classification [63, 9, 66].
@@
Local differential privacy (LDP) is a widely accepted mathematical framework for privacy protection through controlled data perturbation. It uses a parameter $\varepsilon$ (the privacy parameter) to quantify the level of privacy protection. When data is processed through an LDP mechanism before being sent to a classifier, it receives provable privacy guarantees, meaning that any observer including the classifier cannot confidently infer the original sensitive data. Researchers have developed numerous LDP mechanisms for various data domains [17, 31, 55, 54, 37, 26], each offering different perturbation strategies while maintaining formal privacy guarantees.
@@
Data utility is the most crucial metric when implementing LDP mechanisms. While privacy protection is controlled by the privacy parameter $\varepsilon$, data utility measures how well the perturbed data serves its intended application purpose. For simple applications like summation queries, utility can be directly quantified through the theoretical mean squared error (MSE) of LDP mechanisms [55, 54]. For classification tasks, utility corresponds to classification accuracy, which cannot be analytically expressed using MSE. A trivial approach to evaluate classifier utility under an LDP mechanism with a given privacy parameter $\varepsilon$ is to repeatedly perturb the data and measure the proportion of correctly classified instances, as illustrated in Figure 1. While practical, this empirical method has significant limitations: it only applies to specific $\varepsilon$ values and perturbed data instances; changing $\varepsilon$ requires time-consuming re-evaluation, and different instances lead to different results. Meanwhile, it fails to establish relationships between utility and privacy parameters, making it inadequate for systematic comparison of LDP mechanisms. Analytical utility quantification, by contrast, provides valuable guidance for classifier designers and users; however, unlike the MSE for summation queries, we currently lack such a quantification framework for classifiers.
@@
Quantifying the relationship between privacy and classifier utility presents significant analytical challenges. These challenges stem from two aspects: (i) LDP mechanisms inherently introduce perturbations across the entire data domain, potentially causing no utility for classifiers. (ii) Classifiers are often complex or even black-box functions, making it difficult to analyze their behavior under perturbations.
@@
This paper develops a framework to theoretically quantify classifier utility under LDP mechanisms. Our approach addresses the aforementioned challenges through two key insights: (i) LDP mechanisms generate perturbed data that concentrates within a bounded region around the original data with a specific high probability, making extreme perturbations rare. (ii) Within this concentration region, the utility analysis of a classifier can be reformulated as a robustness analysis problem. Robustness refers to how well a classifier maintains correct predictions when inputs are subjected to perturbations. We leverage established robustness analysis techniques to find the maximum permissible perturbation that preserves a classifier’s output, thereby maintaining utility. By combining the concentration analysis of LDP mechanisms with the robustness analysis of classifiers, we establish theoretical data utility quantification in relation to the privacy parameter $\varepsilon$, expressed as:
@@
Given classifier $h$, LDP mechanism $\mathcal{M}_{\varepsilon}$, and raw data $x$: with probability at least $p(\varepsilon,\theta)$, $h$ preserves its correct classification result under $\mathcal{M}_{\varepsilon}(x)$.
@@
This formulation provides a probabilistic guarantee for preserving the utility of classifier $h$ when using mechanism $\mathcal{M}_{\varepsilon}$. The parameter $\theta$ represents a concentration threshold that determines the probability guarantee $p(\varepsilon,\theta)$, which directly depends on the privacy parameter $\varepsilon$. Figure 1 illustrates this intuition. By determining the maximum allowable perturbation $B_{\theta}(x)$ that a classifier $h$ can tolerate, we can derive utility guarantees for $h$ under $\mathcal{M}_{\varepsilon}$ across different privacy levels without needing to repeatedly sample from the mechanism. Furthermore, given the distribution of raw data $x$, our framework also provides average-case and worst-case utility guarantees.
@@
Additionally, we introduce two refinement techniques to refine the utility quantification $p(\varepsilon,\theta)$ for high-dimensional classifiers. First, we extend the widely used robustness notion of “robustness radius” to “robustness hyperrectangle”, enabling a more precise analysis of classifier robustness. Second, we adapt a relaxed privacy notion, Probably Approximately Correct (PAC) privacy [62], as an alternative to $(\varepsilon,\delta)$-LDP. Under this notion, we propose a novel privacy indicator and an extended Gaussian mechanism applicable for any $\varepsilon$. The original Gaussian mechanism from Dwork [17] is only valid for $\varepsilon\in[0,1]$. We improve the proof technique to extend it to $\varepsilon\in[0,\infty)$. These extensions allow the framework to provide more accurate utility quantification and incorporate more privacy-preserving mechanisms.
@@
Applications. The proposed utility quantification framework has direct applications in privacy-preserving classification systems: (i) It enables a comparative analysis of different LDP mechanisms for a given classifier by evaluating their probability guarantees $p(\varepsilon,\theta)$. An LDP mechanism that provides a higher $p(\varepsilon,\theta)$ ensures better utility at the same privacy level. (ii) The framework facilitates the selection of an appropriate privacy parameter $\varepsilon$ to meet specific utility requirements. For example, given a utility threshold $p^{*}$, the framework identifies the $\varepsilon$ that satisfies $p(\varepsilon,\theta)\geq p^{*}$, achieving a precise privacy-utility balance when using the classifier.
@@
  • Quantification framework. To the best of our knowledge, this is the first framework that provides analytical utility quantification for classifiers under LDP mechanisms. This framework bridges the concentration analysis of LDP mechanisms with the robustness analysis of classifiers, enabling systematic evaluation of utility.
@@
  • Comparison and case studies. We conduct case studies on typical classifiers, including logistic regression, random forests, and neural networks, under various LDP mechanisms. The results demonstrate that our theoretical utility quantification aligns closely with empirical observations, particularly in low-dimensional input spaces.
Structure. The main parts of this paper are organized as follows: Section 3 provides an illustrative example of utility quantification for a classifier under the Laplace mechanism. Section 4 introduces the formal framework for utility quantification. Section 5 presents refinement techniques, then Section 7 conducts case studies.
2 Preliminaries
This section formulates the problem and reviews the definition LDP as well as the concept of classifier robustness.
2.1 Problem Formulation
@@
2.2 Local Differential Privacy
Definition 1 ($\varepsilon$-LDP [15]).
A perturbation mechanism $\mathcal{M}:\mathcal{X}\to\mathsf{Range}(\mathcal{M})$ satisfies $\varepsilon$-LDP if, for any two arbitrary inputs $x_{1}$ and $x_{2}$, the probability ratio of producing the same output $\tilde{x}$ is bounded as follows:
@@ 2.2 Local Differential Privacy
If $\mathcal{M}(x)$ is continuous, the probability $\Pr[\cdot]$ is replaced by the probability density function (pdf). Definition 1 quantifies the difficulty of distinguishing between $x_{1}$ and $x_{2}$ based on the output $\tilde{x}$. A smaller privacy parameter $\varepsilon\in[0,+\infty)$ indicates stronger privacy protection. For convenience, we use $\mathcal{M}_{\varepsilon}$ to denote an $\varepsilon$-LDP mechanism $\mathcal{M}$ when $\varepsilon$ is necessary.
Some LDP mechanisms, such as the Laplace and Gaussian mechanisms [17], were originally designed for query functions on raw data. These mechanisms require the sensitivity of the query function to determine the scale of perturbation. The (global) sensitivity of a function $f:\mathcal{X}\to\mathsf{Range}(f)$ is defined as the maximum change in $f$ for any two data:
@@
In the context of LDP, the privacy of the data $x$ itself must be ensured, as the perturbation is applied directly to the data. Here, the query function $f$ is the identity function $f(x)=x$, whose sensitivity is the maximum length of $\mathcal{X}$. For classifiers, the input space $\mathcal{X}$ is typically normalized to $\mathcal{X}=[0,1]$, resulting in a sensitivity of $1$.
2.3 Classifier and Robustness
@@
3 Illustrating the Framework: An Example
@@
Definition 4 (The Laplace mechanism [17]).
For any $x\in[0,1]$, the Laplace mechanism $\mathcal{M}_{\varepsilon}$ is defined as: $\mathcal{M}_{\varepsilon}(x)=x+\xi$, where $\xi\sim Lap(1/\varepsilon)$ are random variables drawn from Laplace distribution with mean 0 and scale $1/\varepsilon$.
In the Laplace mechanism, $\mathcal{M}_{\varepsilon}(x)$ is unbounded due to the Laplace distribution. Since the classifier is typically defined on a bounded domain, such as $[0,1]$, a common practice is to truncate the perturbed data to this range, i.e. $\mathcal{M}_{\varepsilon}(x)$ is truncated to [0,1][0,1]. [^2] This truncation can be treated as a post-processing step by the users that does not leak privacy [17, 57, 25]. We will also examine mechanisms with bounded perturbations in Section 4.1, where no truncation is needed. There are also truncated Laplace mechanisms [11, 23] that directly sample from the truncated Laplace distribution. It is also applicable to our framework; we use the original Laplace mechanism on $[0,1]$ as an example for simplicity.
3.1 Example
@@
Figure 2: Illustration of the example and utility quantification of the classifier under the Laplace mechanism for $x=0.5$.
@@
To quantify the utility of $h(\mathcal{M}_{\varepsilon}(0.5))$ with respect to $\varepsilon$, we analyze two key properties: the concentration of $\mathcal{M}_{\varepsilon}$ and the robustness of $h$.
Concentration analysis. Denote region $B_{\theta}(x)$ as the $\theta$-neighborhood ball of $x$. For instance, $B_{0.3}(0.5)\coloneq\{\tilde{x}:|\tilde{x}-0.5|\leq 0.3\}$. Then the perturbed data $\mathcal{M}_{\varepsilon}(0.5)$ falls into $B_{0.3}(0.5)$ with a probability $p(\varepsilon,\theta=0.3)$. Specifically, denote $F_{\text{Lap}}(\theta)$ as the cumulative distribution function (CDF) of the Laplace distribution used in $\mathcal{M}_{\varepsilon}$. We have:
@@
Robustness analysis. Analyzing the utility of $h(\mathcal{M}_{\varepsilon}(0.5))$ requires analyzing the robustness of $h$ under the input range $B_{\theta}(0.5)$. We can see that $h$ is robust under the input range $B_{0.3}(0.5)=[0.2,0.8]$ according to its definition. Then, we can state that: $h$ preserves the correct classification result under the input range $B_{0.3}(0.5)$.
Utility quantification. Combining the above, we can quantify the utility of $h(\mathcal{M}_{2}(0.5))$: with probability at least $p(2,0.3)=0.46$, $h$ preserves the correct classification result under $\mathcal{M}_{2}(0.5)$. Importantly, this utility quantification can be extended to any $\varepsilon$ using the closed-form expression of $p(\varepsilon,\theta)$.
Application. For this classifier, the theoretical utility quantification $p(\varepsilon,0.3)$ can guide the choice of the privacy parameter $\varepsilon$ in the Laplace mechanism to achieve a desired utility level. Figure 2(b) shows $p(\varepsilon,0.3)$ as a function of $\varepsilon$. For instance, to ensure the classifier preserves the correct result with probability $p(\varepsilon,0.3)\geq 0.8$, privacy parameter $\varepsilon$ should be set to at least $5.2$.
3.2 Further Steps

### v3

@@ Abstract.
local differential privacy, classifier utility, robustness analysis, utility quantification
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
1. Introduction
@@
Figure 1. From empirical data utility to theoretical data utility. This paper provides a theoretical quantification of classifier utility under LDP-perturbed inputs by connecting the concentration analysis of LDP mechanisms with the robustness analysis of classifiers.
@@
Classifiers are functions that map input data to class labels, playing a fundamental role in industrial applications ranging from predictive modeling and data clustering to image recognition (doi:10.1080/01431160600746456; DBLP:journals/information/KowsariMHMBB19; DBLP:journals/csur/JainMF99). When deployed as services, classifiers require users to provide data inputs, which often contain sensitive information such as medical records or financial data, raising significant privacy concerns. While users wish to benefit from classification services, they may be reluctant to reveal their actual data. A common approach to address this concern is data perturbation, e.g. adding noise to numerical data or applying blur filters to images before sending them to the classifier. Data perturbation remains a lightweight and intuitive solution for privacy-preserving classification (DBLP:conf/kdd/ZhangWZ05; DBLP:conf/icdm/ChenL05; DBLP:journals/widm/ZorarpaciO21).
@@
Local differential privacy (LDP) is a widely accepted mathematical framework for privacy protection through controlled data perturbation. It uses a parameter $\varepsilon$ (the privacy parameter) to quantify the level of privacy protection. When data is processed through an LDP mechanism before being sent to a classifier, it receives provable privacy guarantees, meaning that any observer including the classifier cannot confidently infer the original sensitive data. Researchers have developed numerous LDP mechanisms for various data domains (DBLP:journals/fttcs/DworkR14; DBLP:conf/nips/KairouzOV14; DBLP:conf/uss/WangBLJ17; DBLP:conf/icde/WangXYZHSS019; DBLP:conf/sigmod/Li0LLS20; DBLP:conf/uss/Gu0C0020), each offering different perturbation strategies while maintaining formal privacy guarantees.
@@
Data utility is the most crucial metric when implementing LDP mechanisms. While privacy protection is controlled by the privacy parameter $\varepsilon$, data utility measures how well the perturbed data serves its intended application purpose. For simple applications like summation queries, utility can be directly quantified through the theoretical mean squared error (MSE) of LDP mechanisms (DBLP:conf/uss/WangBLJ17; DBLP:conf/icde/WangXYZHSS019). For classification tasks, utility corresponds to classification accuracy, which cannot be analytically expressed using MSE. A trivial approach to evaluate classifier utility under an LDP mechanism with a given privacy parameter $\varepsilon$ is to repeatedly perturb the data and measure the proportion of correctly classified instances, as illustrated in Figure 1. While practical, this empirical method has significant limitations: it only applies to specific $\varepsilon$ values and perturbed data instances; changing $\varepsilon$ requires time-consuming re-evaluation, and different instances lead to different results. Meanwhile, it fails to establish relationships between utility and privacy parameters, making it inadequate for systematic comparison of LDP mechanisms. Analytical utility quantification, by contrast, provides valuable guidance for classifier designers and users; however, unlike the MSE for summation queries, we currently lack such a quantification framework for classifiers.
@@
Quantifying the relationship between privacy and classifier utility presents significant analytical challenges. These challenges stem from two aspects: (i) LDP mechanisms inherently introduce perturbations across the entire data domain, potentially causing zero utility for classifiers. (ii) Classifiers are often complex or even black-box functions, making it difficult to analyze their behavior under perturbed inputs.
@@
This paper develops a framework to theoretically quantify classifier utility under LDP-perturbed inputs. Our approach addresses the aforementioned challenges through two key insights: (i) LDP mechanisms generate perturbed data that concentrates within a bounded region around the original data with a specific high probability, making extreme perturbations rare. (ii) Within this concentration region, the utility analysis of a classifier can be reformulated as a robustness analysis problem. Robustness refers to how well a classifier maintains correct predictions when inputs are subjected to perturbations. We leverage established robustness analysis techniques to find the maximum permissible perturbation that preserves a classifier’s output, thereby maintaining utility. By combining the concentration analysis of LDP mechanisms with the robustness analysis of classifiers, we establish theoretical data utility quantification in relation to the privacy parameter $\varepsilon$, expressed as:
@@
Given classifier $h$, LDP mechanism $\mathcal{M}_{\varepsilon}$, and raw data $x$: with probability at least $p(\varepsilon,\theta)$, $h$ preserves its correct classification under input $\mathcal{M}_{\varepsilon}(x)$.
@@
This formulation provides a probabilistic guarantee for preserving the utility of classifier $h$ when using mechanism $\mathcal{M}_{\varepsilon}$. The parameter $\theta$ represents a concentration threshold that determines the probability guarantee $p(\varepsilon,\theta)$, which directly depends on the privacy parameter $\varepsilon$. Figure 1 illustrates this intuition. By determining the maximum allowable perturbation $B_{\theta}(x)$ that a classifier $h$ can tolerate, we can derive utility guarantees for $h$ under $\mathcal{M}_{\varepsilon}$-perturbed inputs across different privacy parameters $\varepsilon$, without needing to repeatedly sample from the mechanism. Furthermore, if the distribution of raw data $x$ is known, our framework also provides average-case and worst-case utility quantifications.
@@
Additionally, we introduce two refinement techniques to refine the utility quantification $p(\varepsilon,\theta)$ for high-dimensional classifiers. First, we extend the widely used robustness notion of “robustness radius” to “robustness hyperrectangle”, enabling a more precise analysis of classifier robustness. Second, we adapt a relaxed privacy notion, Probably Approximately Correct (PAC) privacy (DBLP:conf/crypto/XiaoD23), as an alternative to $(\varepsilon,\delta)$-LDP. Under this notion, we propose a novel privacy indicator and an extended Gaussian mechanism [^1] It is well known that the Gaussian mechanism cannot satisfy pure (L)DP (DBLP:journals/fttcs/DworkR14); it must rely on a relaxed privacy notion such as $(\varepsilon,\delta)$-LDP or PAC LDP. applicable for any $\varepsilon$. The original Gaussian mechanism from Dwork (DBLP:journals/fttcs/DworkR14) is only valid for $\varepsilon\in[0,1]$. We improve the proof technique to extend it to $\varepsilon\in[0,\infty)$. These extensions allow the framework to provide more accurate utility quantification and incorporate more privacy-preserving mechanisms.
@@
Applications. The proposed utility quantification framework has direct applications in privacy-preserving classification systems: (i) It enables systematic comparison of different LDP mechanisms for a given classifier by evaluating their probability guarantees $p(\varepsilon,\theta)$. For a fixed $\varepsilon$, a mechanism with a larger $p(\varepsilon,\theta)$ provides stronger utility preservation. (ii) It facilitates the selection of an appropriate privacy parameter $\varepsilon$ to meet specific utility requirements. For example, given a utility threshold $p^{*}$, the framework identifies the $\varepsilon$ that satisfies $p(\varepsilon,\theta)\geq p^{*}$, achieving a precise privacy-utility balance when using the classifier.
@@
  • Quantification framework. To the best of our knowledge, this is the first framework that provides analytical utility quantification for classifiers under LDP-perturbed inputs. This framework bridges the concentration analysis of LDP mechanisms with the robustness analysis of classifiers, enabling systematic evaluation of classifier utility.
@@
  • Comparison and case studies. We conduct case studies on typical classifiers, including logistic regression, random forests, and neural networks, under input perturbations by various LDP mechanisms. The results demonstrate that our theoretical utility quantification aligns closely with empirical observations, particularly in low-dimensional input spaces.
Structure. The main parts of this paper are organized as follows: Section 3 provides an illustrative example of utility quantification for a classifier under the Laplace mechanism. Section 4 introduces the formal framework for utility quantification. Section 5 presents refinement techniques, then Section 6 discusses related questions and Section 7 conducts case studies.
2. Preliminaries
This section formulates the problem and reviews the definition of LDP as well as the concept of classifier robustness.
2.1. Problem Formulation
@@
2.2. Local Differential Privacy
Definition 0 ($\varepsilon$-LDP (DBLP:journals/corr/DuchiWJ16)).
A perturbation mechanism $\mathcal{M}:\mathcal{X}\to\mathsf{Range}(\mathcal{M})$ satisfies $\varepsilon$-LDP if, for any two arbitrary $x_{1}$ and $x_{2}$, the probability ratio of producing the same $\tilde{x}$ is bounded as follows:
@@ 2.2. Local Differential Privacy
If $\mathcal{M}(x)$ is continuous, the probability $\Pr[\cdot]$ is replaced by the probability density function (pdf). Definition 1 quantifies the difficulty of distinguishing between $x_{1}$ and $x_{2}$ based on the observed $\tilde{x}$. A smaller privacy parameter $\varepsilon\in[0,+\infty)$ indicates stronger privacy protection. For convenience, we use $\mathcal{M}_{\varepsilon}$ to denote an $\varepsilon$-LDP mechanism $\mathcal{M}$ when $\varepsilon$ is necessary.
Some LDP mechanisms, such as the Laplace and Gaussian mechanisms (DBLP:journals/fttcs/DworkR14), were originally designed for query functions on raw data. These mechanisms require the sensitivity of the query function to determine the scale of perturbation. The (global) sensitivity of a function $f:\mathcal{X}\to\mathsf{Range}(f)$ is defined as the maximum change in $f$ for any two data:
@@
In the context of LDP, the privacy of the data $x$ itself must be ensured, as the perturbation is applied directly to the data. Thus, here the query function $f$ is the identity function $f(x)=x$, whose sensitivity is the diameter of input space $\mathcal{X}$. For classifiers, $\mathcal{X}$ is typically normalized to $\mathcal{X}=[0,1]$, resulting in a sensitivity of $1$.
2.3. Classifier and Robustness
@@
3. Illustrating the Framework: An Example
@@
Definition 0 (The Laplace mechanism (DBLP:journals/fttcs/DworkR14)).
For any $x\in[0,1]$, the Laplace mechanism $\mathcal{M}_{\varepsilon}$ is defined as: $\mathcal{M}_{\varepsilon}(x)=x+\xi$, where $\xi\sim Lap(1/\varepsilon)$ is a random variable drawn from Laplace distribution with mean 0 and scale $1/\varepsilon$.
In the Laplace mechanism, $\mathcal{M}_{\varepsilon}(x)$ is unbounded due to the Laplace distribution. Since the classifier is typically defined on a bounded domain, such as $[0,1]$, a common practice is to truncate the perturbed data to this range, i.e. $\mathcal{M}_{\varepsilon}(x)$ is truncated to [0,1][0,1]. [^3] This truncation can be treated as a post-processing step by the users that does not leak privacy (DBLP:journals/fttcs/DworkR14; DBLP:conf/ndss/0001LLSL20; DBLP:conf/stoc/GhoshRS09). We will also examine mechanisms with bounded perturbations in Section 4.1, where no truncation is needed. There are also truncated Laplace mechanisms (DBLP:journals/jcst/CroftSS22; DBLP:conf/aistats/GengDGK20) that directly sample from the truncated Laplace distribution. It is also applicable to our framework; we use the original Laplace mechanism on $[0,1]$ as an example for simplicity.
3.1. Example
@@
Figure 2. Illustration of the example and utility quantification of the classifier under the Laplace mechanism for $x=0.5$.
@@
To quantify the utility of $h(\mathcal{M}_{\varepsilon}(0.5))$ with respect to $\varepsilon$, we analyze two key properties: the concentration of LDP mechanism $\mathcal{M}_{\varepsilon}$ and the robustness of classifier $h$.
Concentration analysis. Denote region $B_{\theta}(x)$ as the $\theta$-neighborhood ball of $x$. For instance, $B_{0.3}(0.5)\coloneq\{\tilde{x}:|\tilde{x}-0.5|\leq 0.3\}$. Then the perturbed data $\mathcal{M}_{\varepsilon}(0.5)$ falls within $B_{0.3}(0.5)$ with a probability $p(\varepsilon,\theta=0.3)$. Specifically, denote $F_{\text{Lap}}(\theta)$ as the cumulative distribution function (CDF) of the Laplace distribution used in $\mathcal{M}_{\varepsilon}$. We have:
@@
Robustness analysis. Analyzing the utility of $h(\mathcal{M}_{\varepsilon}(0.5))$ requires analyzing the robustness of $h$ under the input range $B_{\theta}(0.5)$. We know that $h$ is robust under input range $B_{0.3}(0.5)=[0.2,0.8]$ according to its definition. Then, we can state that: $h$ preserves the correct classification under input range $B_{0.3}(0.5)$.
Utility quantification. Combining the above, we can quantify the utility of $h(\mathcal{M}_{2}(0.5))$: with probability at least $p(2,0.3)=0.46$, $h$ preserves the correct classification under input $\mathcal{M}_{2}(0.5)$. Importantly, this utility quantification can be extended to any $\varepsilon$ using the closed-form expression of $p(\varepsilon,\theta)$.
Application. For this classifier, the theoretical utility quantification $p(\varepsilon,0.3)$ can guide the choice of the privacy parameter $\varepsilon$ in the Laplace mechanism to achieve a desired utility level. Figure 2(b) shows $p(\varepsilon,0.3)$ as a function of $\varepsilon$. For instance, to ensure the classifier preserves the correct classification with probability $\geq 0.8$, privacy parameter $\varepsilon$ should be set to at least $5.2$.
3.2. Further Steps